Divide rs 300 between Gina and Rina such that Rina gets Rs 100 less than 3 times of what
Gina gets.

Answer by stanbon(75887)

Divide rs 300 between Gina and Rina such that Rina gets Rs 100 less than 3 times of what Gina gets.
------
G + R = 300
R = 3G-100
--------------------
Substitute and solve for "G"::
G + 3G-100 = 300
4G = 400
Gina = 100 rs
----
R = 3G-100 = 300-100 = 200 rs
--------------------
Cheers,
Stan H.